An aqueous solution containing 4% nonvolatile solute exerts a pressure of 1.002 bar at the normal boiling point of the solvent. What is the molar mass of the solute?

Grade:12

1 Answers

Avinash
8 years ago

Vapour pressure of the solution at normal boiling point (p1) = 1.004 bar
Vapour pressure of pure water at normal boiling point
Mass of solute, (w2) = 2 g
Mass of solvent (water), (w1) = 98 g
Molar mass of solvent (water), (M1) = 18 g mol−1
According to Raoult’s law,
pi0-p/pi0=w2*m1/(w1*m2)
1.013-1.004/1.013=2*18/98*m2
= 41.35 g mol−1
Hence, the molar mass of the solute is 41.35 g mol−1.